# Advancing Structure Prediction of Biomolecular Interaction via Contact-Guided Sampling with HelixFold-S1

### v1

@@ Abstract
Biomolecular structure prediction is essential to molecular biology, yet accurately predicting the structures of complexes remains challenging, especially when co-evolutionary signals are absent. While recent methods have improved prediction accuracy through extensive sampling, aimless sampling often provides diminishing returns due to limited conformational diversity. Here, we introduce HelixFold-S1, a contact-guided sampling strategy that improves structural accuracy. Rather than relying on indiscriminate sampling, HelixFold-S1 predicts contact probabilities between molecular entities and uses these predictions to prioritize sampling of likely binding sites and modes. This targeted approach generates a diverse set of structural candidates, enhancing the likelihood of identifying accurate conformations. We demonstrate that HelixFold-S1 consistently outperforms baseline sampling strategies across a range of biomolecular interactions, including protein-antibody, protein–protein, protein–ligand, protein-RNA, and protein–DNA interfaces. Furthermore, the predicted contact probabilities serve as a reliable indicator of structural difficulty, guiding the allocation of sampling resources. These results highlight the potential of targeted sampling strategies to advance the structural modeling of complex biomolecular interactions during inference.
@@
Biomolecular interaction, Structure Prediction, Test-time scaling, Sampling Strategy
@@ Introduction
Biomolecular structure prediction is a cornerstone of computational biology, underpinning advances in drug discovery, protein engineering, and our broader understanding of molecular interactions. Recent breakthroughs by deep learning-based methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], represented by AlphaFold [1, 2, 3] and RoseTTAFold series [4, 5], have transformed the field by exploiting co-evolutionary signals and leveraging deep neural networks to infer complex structural relationships.
Despite these advances, predicting the structures of biomolecular complexes remains challenging, particularly when co-evolutionary signals are weak or unavailable. This is especially true for antigen–antibody pairs, where antibody diversity and the lack of co-evolution with antigens hinder accurate interface prediction. These limitations are further compounded by the difficulty of achieving high accuracy from a single predicted structure. While the AlphaFold series has sought to improve precision through model ensembling (e.g., five or twenty-five ensembles), the limited number of generated structures constrains comprehensive exploration of potential binding sites and interaction modes, ultimately restricting improvements in accuracy. To address this, recent studies [12, 13, 14, 15] have explored large-scale sampling strategies aimed at increasing conformational diversity by generating thousands of candidate structures. For example, AFSample [12] and AFSample2 [13] apply dropout and random MSA column masking during inference to produce diverse conformations, while AlphaFold3 [3] demonstrates that extensive sampling can significantly enhance predictions of antigen–antibody interfaces. These approaches share conceptual similarities with test-time scaling in large language models (LLMs) [16, 17, 18, 19, 20, 21], where greater computational resources are allocated during inference to improve output quality. Nevertheless, the structural variability of samples produced by folding models remains relatively low, leading to diminishing gains from repeated sampling. Accordingly, increasing the number of predictions alone provides only limited benefits in terms of accuracy. Consequently, merely increasing the number of predictions provides marginal improvements in accuracy. This highlights the necessity for more intelligent, resource-efficient sampling strategies that can prioritize the exploration of structurally informative regions and achieve higher accuracy with reduced computational overhead.
To address the inefficiencies of traditional large-scale sampling, we propose HelixFold-S1, which introduces a contact-guided sampling strategy to more efficiently explore the conformational space and improve structural prediction accuracy. Rather than exploring the conformational space randomly or exhaustively, HelixFold-S1 estimates a probability distribution over potential intermolecular contacts, prioritizing regions with a high likelihood of interaction. These predicted contact regions are used as spatial constraints to guide the model’s structural predictions, directing computational resources toward areas that provide the most structural insight. The generated conformations are then evaluated and ranked based on the associated confidence scores. This targeted approach enhances the structural diversity of the predictions, expanding the search for plausible binding modes and increasing the likelihood of identifying accurate structures.
We evaluate HelixFold-S1 across a range of biomolecular interaction scenarios, including protein–protein, protein–antibody, protein–ligand, protein–RNA, and protein–DNA interfaces. Compared to traditional sampling strategies, the contact-guided sampling approach employed by HelixFold-S1 demonstrates a marked improvement in prediction accuracy as the sampling process is increased. This advantage is particularly evident in more challenging scenarios, such as protein–antibody interactions with limited co-evolutionary signals and protein–ligand interfaces involving ligands absent from the training set. Moreover, the contact probabilities predicted by HelixFold-S1 exhibit a clear correlation with the difficulty of structural prediction. Specifically, targets with lower predicted contact probabilities tend to yield lower prediction accuracy, indicating greater structural complexity, while higher contact probabilities correspond to improved accuracy. For targets with intermediate contact probabilities, HelixFold-S1 shows a pronounced potential for improvement with increased sampling. Additionally, the conformational ensembles generated by HelixFold-S1 are more diverse, encompassing a broader spectrum of potential contact regions. This increased diversity enhances the robustness of the structural predictions and facilitates the identification of near-native conformations. HF-S1 has been deployed on the PaddleHelix platform and is available for online use.
@@ Results
Framework of HelixFold-S1
Figure 1: Overall framework of HelixFold-S1. a, Inference pipeline of HF-S1: a three-stage process for contact-guided structure prediction. Stage 1: a contact prediction module generates inter-chain contact probabilities. Stage 2: an algorithm samples diverse contact subsets. Stage 3: selected contacts guide structure generation and ranking via the contact conditioning module. b, Network architecture of HF-S1. Input sequences, MSA/template features, and optional contact constraints are passed to the Input Embedder, with contact constraints processed by the contact conditioning module. The Pairformer models these inputs, followed by a diffusion module predicting the 3D structure. A contact prediction module estimates inter-token contact probabilities. During training, the contact conditioning and prediction modules are activated mutually exclusively.
HelixFold-S1 (HF-S1) is a sampling-optimized variant of HelixFold3 (HF3) [9], an open-source biomolecular structure prediction model built with PaddlePaddle [22] (Fig. 1). HF3 reproduces the core capabilities of AlphaFold3 (AF3) [3], supporting a broad range of biomolecules, including proteins, nucleic acids, and small molecules.
@@
To enhance inference efficiency and structural diversity, HF-S1 introduces two contact-centric modules: the Contact Prediction Module and the Contact Conditioning Module (Fig. 1b). A contact is defined as existing between two tokens if any pair of atoms from the corresponding tokens lies within 5Å. The Contact Prediction Module computes a contact probability matrix across all token pairs using the pair representations produced by the Pairformer, a specialized attention-based module designed to capture long-range pairwise token dependencies. The Contact Conditioning Module, incorporated into the Input Embedder, enables the model to learn the contact constraints and generate structures based on given contact constraints. These two modules are activated in a mutually exclusive manner during training, creating a multi-task framework that alternates between contact probability estimation and contact-conditioned structure generation. This allows the model to jointly learn to infer inter-chain interactions and effectively apply them during inference.
The inference pipeline of HF-S1 consists of three stages (Fig. 1a). Contact Prediction: HF-S1 first computes a contact probability matrix over all token pairs, estimating the likelihood that any two tokens are in spatial proximity (i.e., any atom pair within 5Å). However, only the entries corresponding to inter-chain token pairs are considered valid, as the model is designed to focus on contacts between different molecular chains. This filtered matrix defines a probabilistic landscape of inter-chain interactions and guides the subsequent sampling process. Contact Sampling: To efficiently explore the structural landscape, HF-S1 employs a greedy sampling strategy that selects contacts in descending order of their predicted probabilities. Importantly, to reduce sampling redundancy, once a conformation is predicted using a particular contact as a constraint, any additional contacts that are also satisfied within the resulting conformation will be excluded from subsequent sampling. We call this strategy as redundant contact pruning. Contact-Guided Structure Prediction and Ranking: For each sampled contact, the model switches to contact conditioning mode to generate a structure that satisfies the contact constraint. All resulting structures are then ranked based on model confidence scores, and the top-ranked prediction is selected as the final output.
@@
Figure 2: Improved structural accuracy across complex types. a, Representative example illustrating HF-S1’s sampling strategy. The heatmap depicts predicted inter-chain contact probabilities between residues of protein A and protein B. Surface residues are colored according to their highest predicted contact probability with the partner chain. b, Structural accuracy of HF-S1 evaluated on benchmark complexes collected from the RCSB PDB between January 1, 2022 and December 31, 2024. Accuracy improves with increased sampling, and HF-S1 consistently achieves high performance across a range of interface types, including protein–antibody (n=221), protein–protein (n=199), protein–ligand (n=194), protein–RNA (n=94), and protein–DNA (n=252). c, Comparison of HF-S1 to various baselines on protein-antibody complexes released in year 2024 (n=95) with sampling number of 80. d, Accuracy on protein-antibody complexes of year 2024 (n=95) changes with the sampling number.
@@
To systematically evaluate the performance of HF-S1 across a range of biologically relevant complex types, we constructed a test set comprising protein–antibody (n=221), protein–protein (n=199), protein–ligand (n=194), protein–RNA (n=94), and protein–DNA (n=252) complexes, collected between 2022.01.01 and 2024.12.31 from the RCSB PDB [23]. To minimize overlap with the training data, all test samples were selected to have low sequence identity to the training set. Sequences were clustered by similarity [24], and one representative per cluster was randomly chosen to ensure diversity and reduce redundancy. For protein–ligand complexes, any ligands that appeared in the training data were further excluded.
We benchmarked HF-S1 against HF3 [9], an open-source implementation of AF3 that achieves accuracy comparable to the original model. For both models, multiple structural conformations were sampled per target. Prediction performance was evaluated using two selection strategies: score-based selection, in which the top-ranked conformation was chosen based on the model’s predicted confidence scores, and oracle selection, in which the most accurate conformation was retrospectively selected based on the experimental structure. For protein–protein complexes (including protein–antibody), performance was assessed using average DockQ scores [25]. For protein–ligand complexes, accuracy was measured by the proportion of predictions with a pocket-aligned ligand root mean squared deviation (RMSD) below 2Å. For protein–nucleic acid complexes (including protein–RNA and protein–DNA), interface accuracy was evaluated using interface LDDT (iLDDT) [26].
We first illustrate the model’s sampling strategy, we present a representative example of a complex structure predicted by HF-S1 (Fig. 2a). The model begins by generating an inter-chain contact probability matrix—visualized as a heatmap—that quantifies the likelihood of contact between each token pair from protein A and protein B. To highlight potential interaction regions, the surface residues of each protein are colored based on the highest contact probability in their respective rows or columns. This visualization effectively identifies residues most likely to participate in inter-chain interactions. Using this contact map as guidance, HF-S1 samples multiple plausible structural conformations, capturing diverse binding orientations between the two proteins.
Across all complex types (Fig. 2b), both HF3 and HF-S1 exhibited improved accuracy as the number of sampled conformations increased, underscoring the significance of test-time scaling in biological systems. The widening performance gap under oracle selection between HF-S1 and HF3 with additional conformations highlights the advantages of the advanced contact-guided sampling strategies implemented in HF-S1, which effectively enhance conformational diversity and structural accuracy. Notably, HF-S1 demonstrated particularly substantial improvements for protein–antibody interfaces compared with general protein–protein interfaces. HF-S1 yields greater performance gains over HF3 on protein–antibody complexes than on general protein–protein interactions (Fig. 2c). This is likely due to the absence of co-evolutionary signals in antibody–antigen interactions, which results from the highly diverse and individually generated nature of antibodies and the lack of long-term evolutionary coupling between binding partners. For protein–ligand complexes, the intrinsic flexibility and diverse binding modes of ligands pose additional challenges for traditional methods. In these scenarios, HelixFold-S1’s enhanced sampling mitigates the absence of strong evolutionary constraints, leading to improved prediction accuracy. Furthermore, HF-S1 exhibited greater improvements under the oracle selection strategy than under score-based selection, indicating that the current scoring function may not fully capture the most accurate conformations. This suggests that further optimization of the score-based selection process could help better exploit the benefits of enhanced conformational sampling.
We further benchmark HF-S1 against several representative baselines on protein–antibody complexes (Fig. 2c and 2d), including two state-of-the-art AF3-like structure prediction models, Chai-1 [27] and Boltz-2 [28], as well as the advanced sampling algorithm AFsample [12]. Among them, Chai-1 and Boltz-2 reflect the current leading performance in biomolecular structure prediction, while AFsample represents a strong post-hoc sampling enhancement method. To assess the impact of sampling strategies independently of model backbone quality, we construct a hybrid baseline, HF3 w/ AFsample, by replacing HF3’s original sampling method with that of AFsample. This enables a direct comparison of sampling effectiveness across methods. To ensure fair evaluation and avoid data leakage—particularly considering that Boltz-2 was trained on PDB entries up to June 1, 2023—we curate a test subset of 95 protein–antibody complexes released after January 1, 2024. All models are evaluated under the same condition of 80 samples per target. Using DockQ ¿ 0.23 as the success criterion (Fig. 2d), HF3 and Boltz-2 achieve strong baseline accuracy, both outperforming Chai-1. However, their performance shows limited improvement as the number of samples increases, indicating that their native sampling strategies are less effective at exploring alternative conformations. Chai-1, despite being a leading AF3-like model, shows virtually no gain from additional sampling, suggesting minimal responsiveness to increased sample size. By contrast, HF3 w/ AFsample yields noticeable accuracy gains with more samples, confirming the advantage of more sophisticated sampling techniques. Most notably, HF-S1 exhibits significantly faster and more substantial improvements than all baselines, highlighting its superior ability to efficiently explore the structural landscape and refine predictions through sampling.
Diversity and Quality of the Sampled Conformations
Figure 3: Diversity and quality of the sampled conformations. a, Distribution of structural accuracy across all sampled conformations at each sampling step, averaged over multiple datasets. We report the highest, lowest, and 25th-percentile precision scores to characterize both the range and quality of sampled structures. b, Representative examples showing that HF-S1 produces more diverse and higher-quality conformations than HF3, as reflected in the broader distributions of confidence and precision.
HF-S1 outperforms HF3 primarily due to its contact-guided sampling strategy, which stands in contrast to the unguided (random) sampling employed by HF3. We hypothesize that the key advantage of HF-S1 lies in its ability to generate a broader and higher-quality distribution of sampled conformations. Specifically, increased sampling diversity may raise the likelihood of capturing high-precision structures during the inference process.
To evaluate this hypothesis, we analyzed the distribution of structural accuracy across all sampled conformations at each sampling step. For each target and sampling num $s$, we sorted the $s$ generated conformations by their precision scores and tracked three statistics: the highest precision, the lowest precision, and the 25th-percentile precision (i.e., the precision of the conformation ranked at the top 25%). The averaged results across multiple datasets are presented in Fig. 3a. This analysis reveals two key patterns. First, as the number of samples increases, both HF3 and HF-S1 exhibit a broader spread in structural precision—indicating that later-stage sampling leads to both better and worse structures. Second, and more importantly, HF-S1 consistently produces a wider precision range compared to HF3, confirming that the contact-guided approach enhances structural diversity. Furthermore, the 25th-percentile precision values for HF-S1 are consistently higher than those of HF3, indicating that HF-S1 tends to generate a larger fraction of high-quality conformations.
We further illustrate the differences between HF3 and HF-S1’s sampling strategies using two representative examples (Fig. 3b). The first example is the complex structure of fungal $\beta$-1,3-glucanosyltransferases (Gel4) and Nb4 nanobody, where the nanobody binds to a dissimilar CBM43 domain of Gel4 across fungal species (PDB ID: 8pe1) [29]. The second example is a protein-RNA heterodimer, illustrating the interaction between the MD-4HB protein and helix 44 RNA in the yeast ribosome (PDB ID: 7x34) [30]. For each case, we examined the distribution of model-predicted confidence scores and interface-level precision metrics across the sampled conformations. HF-S1 consistently generates a broader spectrum of structures, spanning a wide range of confidence levels and precision values—including a notable fraction of high-accuracy predictions. In contrast, HF3 tends to sample conformations within a narrower confidence and precision range, indicating more limited structural diversity and fewer high-quality candidates.
Predicted Contact Probability
Figure 4: Predicted Contact Probability. a, Accuracy of the contact probability predicted by HF-S1. b, Contact satisfaction rate, i.e., fraction of structures predicted by HF-S1 satisfying contact constraints. c, Correlation between the target-level contact probability (maximum value in the predicted contact probability matrix) and the accuracy of the predicted structure. d, Improvement from multiple sampling by HF-S1 across different target contact probability groups under oracle selection and score-based selection.
@@
The contact probability matrix predicted by HF-S1 serves a dual role: it not only informs the structural sampling strategy but also provides insight into the intrinsic difficulty of the structure prediction task, as well as the potential benefits of additional sampling.
We first evaluate the accuracy of the contact probability matrix outputted by HF-S1. For each target, token pairs that form true contacts, as defined by experimental structures, are treated as positive examples, while all others are considered negative. Using this approach, we calculate the area under the precision–recall curve (AUPRC) between the predicted and ground-truth contact maps, reporting the average values across all targets. As a baseline, for each target, we generate a posterior contact probability matrix for HF3 based on the sampled conformations. The matrix element for each token pair is defined as the inverse of the minimum distance observed across the sampled conformations. In comparison to this baseline, the contact probabilities predicted by HF-S1 consistently show strong accuracy across a variety of molecular types (Fig. 4a). Among these categories, protein–antibody complexes exhibit the lowest AUPRC, further confirming the difficulty of this prediction scenario. Interestingly, despite the lack of coevolutionary signals in protein–ligand binding, protein–ligand complexes achieve the highest AUPRC. This may be attributed to the fact that binding sites in protein–ligand interactions are often structurally conserved and spatially well-defined. Many such interfaces exhibit clear geometric features—such as hydrophobic pockets and polar residue arrangements—that help the model accurately localize the binding site.
We next assessed whether the conformations predicted by HF-S1 satisfy the input contact constraints. Specifically, we evaluated the contact satisfaction rate, defined as the fraction of predicted structures in which the specified contacts are realized (Fig. 4b). Across most test cases, HF-S1 consistently adhered to the provided constraints, achieving satisfaction rates above 70% across various types of targets, demonstrating the model’s ability to effectively incorporate such priors into structure prediction.
@@
It is of interest to examine whether predicted contact probabilities can serve as a proxy for the intrinsic difficulty of a target. To this end, we analyzed the relationship between the target-level contact probability—defined as the highest value in the predicted contact probability matrix for each target—and the prediction precision of HF-S1 under single sampling (Fig. 4c). Across all benchmark datasets, we observed a strong correlation: targets with lower contact probabilities tend to exhibit lower structural accuracy, whereas those with higher values are generally predicted more precisely. This suggests that contact probability estimates may reflect the inherent difficulty of the prediction task. When grouped by complex type, protein–protein targets typically show higher target-level contact probabilities, while protein–antibody complexes cluster in the lower range, indicating greater structural uncertainty. Protein–ligand complexes consistently exhibit high contact probabilities, suggesting that HF-S1 can often localize ligand binding sites with high confidence. In contrast, protein–RNA and protein–DNA complexes display a broader distribution, reflecting greater variability in prediction difficulty across these categories.

### v2

# 1]PaddleHelix Team, Baidu Inc

@@ Abstract
Generating large ensembles of candidate conformations is standard for improving biomolecular structure prediction. Yet aimless sampling is inefficient and costly, producing many redundant conformations with limited diversity, so additional computation often yields little improvement. Here, we present HelixFold-S1, a guided planning approach that strategically targets the most informative regions of conformational space to produce accurate conformations. For each biomolecule, predicted inter-chain contact probabilities serve as a blueprint of the conformational space, guiding computational effort toward higher-probability, low-redundancy contacts that constrain structure generation. Across diverse biomolecular benchmarks, HelixFold-S1 achieves markedly higher structural accuracy than traditional unguided methods while reducing sampling requirements by an order of magnitude. Predicted contact probabilities also provide a rough indicator of prediction difficulty and sampling utility. These results demonstrate that guided planning reshapes conformational exploration and enables more efficient and accurate structural inference.
@@
Biomolecular structure prediction, Conformation space exploration, Planning-driven, Computational efficiency, Resource-aware modeling
@@ Introduction
Biomolecular structure prediction lies at the heart of computational biology, enabling progress in drug discovery, protein engineering, and the study of molecular interactions. Deep learning has transformed this field, with recent breakthroughs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] exemplified by the AlphaFold [1, 2, 3] and RoseTTAFold series [4, 5]. Despite these advances, accurate prediction of biomolecular complexes remains challenging.
Improving the accuracy of a single prediction has proved difficult, and generating large ensembles of candidate conformations has therefore emerged as a simple yet powerful way to boost performance. The AlphaFold series improves precision by model ensembling by combining predictions from five or twenty-five networks, and AlphaFold3 [3] further shows that extensive sampling can markedly improve protein–antibody interface modelling. Recent studies [12, 13, 14, 15] have explored large-scale sampling, generating thousands of candidate structures to expand conformational diversity. Approaches such as AFSample [12] and AFSample2 [13] introduce stochasticity during inference, using dropout and random template masking to generate alternative conformations, delivering notable gains in predictive accuracy. Other strategies [16, 17, 18, 19] diversify predictions by repeatedly sub-sampling the input multiple sequence alignments (MSAs), offering complementary routes to probe alternative structural states.
Figure 1: Overall framework of HelixFold-S1. a, Aimless exploration in conventional conformation sampling, often concentrated within limited regions of conformational space. b, Guided exploration in planning-based sampling, directing sampling toward more informative conformations. c, d, Two-stage inference of HelixFold-S1: c, stage 1 predicts a blueprint of the conformational space; d, stage 2 generates conformations based on this blueprint. e, Network architecture of HF-S1, which incorporates a contact conditioning module and a contact prediction module into the HelixFold3’s model architecture. During training, the two modules are activated in a mutually exclusive manner.
Even though generating large ensembles of candidate structures can improve predictive accuracy, it comes at substantial computational cost. For instance, performing a thousand samplings may require an entire GPU day. The main reason such massive sampling is needed is that current sampling methods are largely unguided, leading to aimless exploration of conformational space. As a result, the produced conformations often cluster around a single peak of the conformational landscape (Fig. 1a), yielding only a limited set of highly similar structures. Consequently, most sampled conformations provide little additional information, and overall sampling efficiency remains low. Moreover, it is often unclear whether additional sampling will meaningfully improve predictions, since simpler targets benefit less from extensive ensembling, rendering much of the computation redundant. These challenges highlight the need for planning-driven strategies that can first illustrate the most potentially informative regions, and then intelligently navigate the conformational space to focus computational effort where it matters most, thereby achieving higher accuracy at far lower computational cost.
Here, we introduce HelixFold-S1, a guided sampling strategy that explores conformational space with exceptional efficiency and precision. Instead of relying on random, unguided sampling, HelixFold-S1 strategically targets the most informative regions of the conformational space, with the aim of sampling across multiple higher-probability regions (Fig. 1b). For each biomolecule, we first predict inter-chain inter-residue contact probabilities based on pairwise representations. Interpreting these distributions as a coarse-grained structural blueprint of the conformational space provides a reduced representation that highlights where meaningful interaction patterns are most likely to occur (Fig. 1c). This structural map then guides the sampling, with computational effort deliberately concentrated in regions associated with higher-probability, low-redundancy contacts that act as spatial constraints during structure generation. The resulting conformations are evaluated and ranked by their confidence scores, enabling both accurate and resource-efficient modeling of complex biomolecular assemblies (Fig. 1d).
We evaluated HelixFold-S1 across diverse biomolecular interfaces, including protein–protein, protein–antibody, protein–ligand, protein–RNA, and protein–DNA. Compared with conventional unguided sampling, HelixFold-S1 substantially improves prediction accuracy at equivalent sampling effort, with the largest gains observed for challenging protein–antibody complexes. Remarkably, it achieves comparable accuracy to traditional methods using an order of magnitude fewer sampling steps. The approach is generalizable to other folding models to enhance structural precision. Predicted inter-chain contact probabilities correlate strongly with structural difficulty and provide a practical guide for allocating sampling effort across targets. Guided sampling also improves exploration of the conformational space, producing more diverse and higher-quality ensembles. HelixFold-S1 is deployed on the PaddleHelix platform [^1] https://paddlehelix.baidu.com/app/all/helixfold3/forecast and is available for online use.
@@ Results
Guided Sampling Strategy and Architecture of HelixFold-S1
The inference pipeline of HF-S1 comprises two stages. In the first stage, HF-S1 predicts a coarse blueprint of the conformational space (Fig. 1c) to identify the most critical inter-chain interaction sites. To this end, an additional Contact Prediction Module is employed to estimate inter-chain inter-token contact probabilities based on the pair representation, representing the likelihood that any two tokens are in spatial proximity, i.e., any atom pair lies within 5Å. This contact probability map provides a reduced, coarse-grained representation of the high-dimensional conformational space, serving as a structural blueprint to guide subsequent conformation sampling. In the second stage, conformations are generated guided by the predicted blueprint. Contacts are prioritized according to their predicted probabilities and sequentially selected from the blueprint. Each selected contact is introduced as an additional constraint via the Contact Conditioning Module, which processes these constraints and generates structures that attempt to satisfy them. By focusing on contacts with higher predicted probabilities, the model is more likely to produce multiple high-quality conformations that capture the most informative interactions. To reduce sampling redundancy, we employ a strategy called redundancy contact pruning (RCP). Under this approach, once a conformation is generated using a particular contact, any contacts that are already satisfied within the resulting structure are excluded from further selection. All generated conformations are subsequently ranked according to model confidence scores, with the top-ranked prediction designated as the final output.
@@
The architecture of HelixFold-S1 (HF-S1) builds upon that of HelixFold3 (HF3) [9], which reproduces the biomolecular structure prediction capabilities of AlphaFold3 (AF3) [3], and further extends its capacity to more intelligently explore conformational space through the addition of two contact-centric modules: the Contact Prediction Module (CPM) and the Contact Conditioning Module (CCM), as depicted in Fig. 1e. A contact is considered to exist between two tokens if any pair of atoms from the corresponding tokens lies within 5Å. The Contact Prediction Module computes a contact probability matrix across all token pairs using the pair representations produced by the Pairformer, a specialized attention-based module designed to capture long-range pairwise token dependencies. The Contact Conditioning Module, incorporated into the Input Embedder, enables the model to learn contact constraints and generate structures conditioned on these constraints. During training, these two modules are activated in a mutually exclusive manner, forming a multi-task framework that alternates between contact probability estimation and contact-conditioned structure generation. This design allows the model to jointly learn inter-chain interactions and effectively leverage them during inference. During inference, the two modules are employed sequentially: in the first stage (Fig. 1c), the Contact Prediction Module estimates the overall contact landscape, and in the second stage (Fig. 1d), the Contact Conditioning Module generates conformations guided by the predicted contacts.
@@
Figure 2: Structural performance and module evaluation of HelixFold-S1 across multiple complex types. a, Top-5 structural precision among 25 sampled conformations for various folding models. Benchmark complexes were collected from the RCSB PDB between January 1, 2022, and December 31, 2024, including protein–antibody (n=221), protein–protein (n=198), protein–ligand (n=238), protein–RNA (n=177), and protein–DNA (n=254) complexes. b, Relative improvements of HelixFold-S1 over HelixFold3 across different complex types in Top-1, Top-5, and Top-25 precision. c, Comparison of HF-S1 with various baseline models on protein–antibody complexes released in 2024 (n=74), showing that extensive sampling progressively improves structural accuracy. Some methods were sampled 1000 times, while others used 100 samples. d, Performance comparison between HF3 w/ AFSample and HF-S1 across different antibody types on the same 2024 protein–antibody complexes (n=74), using 1000 samples. e, Performance of the contact prediction module in HF-S1, evaluated using the area under the precision–recall curve (AUPRC). f, Effectiveness of the contact conditioning module, measured as the fraction of predicted structures satisfying contact constraints. g, Example structures predicted by HelixFold-S1 using two contact constraints (PDB ID: 8ozb).
@@
To systematically evaluate the performance of HF-S1 across a range of biologically relevant complex types, we constructed a test set comprising protein–antibody (n=221), protein–protein (n=198), protein–ligand (n=238), protein–RNA (n=177), and protein–DNA (n=254) complexes, collected between 2022.01.01 and 2024.12.31 from the RCSB PDB [20]. To minimize overlap with the training data, all test samples were selected to have low sequence identity to the training set. Sequences were clustered by similarity [21], and one representative per cluster was randomly chosen to ensure diversity and reduce redundancy. For protein–ligand complexes, any ligands that appeared in the training data were further excluded.
We compared AF3-like folding models [^2] AlphaFold3 is not available for commercial use and thus could not be tested. In addition, part of our benchmark overlaps with the Boltz-2 [22] training set, and Boltz-2 was therefore excluded., namely HF3 [9], Protenix [23], and Chai-1 [24], under different sampling strategies. For HF3, we examined two sampling-augmented variants: one following the AFSample strategy [12] (HF3 w/ AFSample) and another implementing our guided planning approach (HF-S1). We further transferred HF-S1–predicted contacts into Protenix and Chai-1, yielding Protenix w/ S1 and Chai-1 w/ S1, to assess their effect on alternative folding models. To enable parallel evaluation, Protenix w/ S1 and Chai-1 w/ S1 were run without redundancy contact pruning (RCP), which requires sequential sampling, whereas HF-S1 fully applies RCP during sampling. Because HF-S1 may employ a distinct definition of contacts from those folding models, the baseline models may not reach their best performance under this evaluation. Moreover, Chai-1 currently supports contact constraints only for protein–protein complexes, and results are therefore reported for protein–antibody and protein–protein systems only. For all methods, we generated 25 predicted structures per target and ranked them using each method’s confidence score.
The precision of the Top-5 ranked structures is shown in Fig. 2a. HF-S1 consistently outperforms both HF3 and HF3 w/ AFSample across all complex types, with the magnitude of improvement highest for protein–antibody complexes, followed by protein–protein and protein–RNA/DNA interfaces, and smallest for protein–ligand binding sites. The largest gains are observed for protein–antibody complexes, where HF-S1 achieves over a 55% improvement relative to HF3 and 33% relative to HF3 w/ AFSample. This likely reflects the structural diversity of antigen-binding sites, which makes conventional sampling less effective at capturing native-like geometries, whereas HF-S1’s guided sampling strategy increases the likelihood of generating correct interfacial contacts. For protein–protein and protein–RNA/DNA interfaces, HF-S1 provides moderate but clear improvements. By enforcing constraints on individual interfacial contacts, guided sampling helps capture correct local interactions, although single-contact constraints alone are not sufficient to fully determine the global geometry of these larger, more complex interfaces. Finally, protein–ligand binding sites are highly constrained, so while HF-S1 can still improve predictions, the relative gains are smaller compared with other complex types. Moreover, across different folding models, including HF3, Protenix, and Chai-1, incorporating the guided sampling strategy of HF-S1 leads to consistent performance gains, underscoring its broad effectiveness. In Fig. 2b, we further demonstrate that the performance advantage of HF-S1 over HF3 is consistently observed across different Top-K precisions of the sampled conformations. The improvements are pronounced for Top-1, Top-5, and Top-25, underscoring HF-S1’s robustness across multiple ranking thresholds, although the Top-1 gains exhibit greater variability, potentially reflecting sensitivity to the selection of conformations based on the confidence score.
Extensive sampling is particularly effective for protein–antibody interfaces [23], and HF-S1 shows the largest gains on these targets. To systematically evaluate the benefit of extensive sampling on these interfaces, we curated a dataset of 74 complexes released after 2024 and with residues less than 800 for efficiency, ensuring no overlap with Boltz-2’s training data [22]. Sampling was conducted with top-performing methods evaluated 1,000 times per complex, while others were sampled 100 times per complex. As shown in Fig. 2c, Top-1 precision on protein–antibody interfaces generally improves with increased sample size across all methods, particularly for those incorporating advanced strategies such as HF-S1 or AFsample. Importantly, HF-S1 achieves comparable precision to HF3 with AFsample using only 10 samples. This indicates that 10 HF-S1 samples are sufficient to reach the precision level that HF3 with AFsample requires 1,000 samples to achieve, representing just 1% of the computational cost. To further assess performance across different antibody types, Fig. 2d presents Top-1 precision for HF-S1 and HF3 with AFsample across canonical antibodies, nanobodies, and scFv based on the full 1,000-sample evaluation. HF-S1 demonstrates consistent improvements in Top-1 precision over HF3 w/ AFsample across all categories.
The Contact Prediction Module (CPM) and Contact Conditioning Module (CCM) are two key additions in HF-S1 built upon the HF3 architecture, and the accuracy of HF-S1 depends on both. To evaluate the CPM, we assessed the accuracy of its predicted contact probability matrices. For each target, token pairs corresponding to true contacts in experimental structures were treated as positives, and all others as negatives. We then computed the area under the precision–recall curve (AUPRC) between predicted and ground-truth contact maps, averaging across all targets. As a baseline, we derived a posterior contact probability matrix for HF3 by aggregating distances from sampled conformations, with each entry defined as the inverse of the minimum inter-residue distance. Compared with HF3, HF-S1 consistently produced more accurate contact probabilities across diverse molecular types (Fig. 2e). Protein–antibody complexes showed the lowest AUPRC, reflecting the difficulty of this prediction scenario, whereas protein–ligand complexes achieved the highest AUPRC, indicating a relatively simpler task. This observation is consistent with the results in Fig. 2a, where protein–antibody interfaces benefit most from advanced sampling strategies, while protein–ligand interfaces show smaller gains under the same approach. To examine the CCM, we evaluated whether HF-S1–predicted conformations adhered to specified contact constraints. We defined the contact satisfaction rate as the fraction of predicted structures in which the given contacts were realized. Across most test cases, HF-S1 achieved satisfaction rates above 70% (Fig. 2f), demonstrating its ability to effectively incorporate contact priors into structure prediction.
Fig. 2g illustrates a representative case. Two inter-chain residue pairs (R43–N26 and V103–L28) were selected from HF-S1’s predicted contact map and applied as spatial restraints during folding, yielding two distinct binding conformations that both satisfied the defined contact distances (R43–N26 $\approx$ 3.1 Å; V103–L28 $\approx$ 4.2 Å).
Contact Probability as an Indicator of Prediction Difficulty and Sampling Utility
Figure 3: Contact probability as an indicator of prediction difficulty and sampling utility. a, Correlation between target-level contact probability (maximum value in the predicted matrix) and structural precision across interface types, evaluated using the Top-5 of 25 sampled conformations. b, Distribution of predicted contact probabilities across interface types, constructed from all benchmark targets in each category. c, Structural improvement from single to extensive sampling by HF-S1 across contact probability groups, measured using the best-performing conformations, showing larger gains for lower-probability interfaces. d, Relative Top-5 precision (of 25 sampled conformations) of HF-S1 compared with HF3 across contact probability levels, indicating stronger benefits on lower-probability targets. e, Predicted contact probabilities reflect residue physicochemical preferences.
@@
The contact probability matrix predicted by HF-S1 serves a dual role: it not only informs the structural sampling strategy but also provides insight into the intrinsic difficulty of the structure prediction task, as well as the potential benefits of multiple sampling.
@@
We first show that predicted contact probabilities can serve as a proxy for estimating target difficulty. Specifically, we defined the target-level contact probability as the maximum value in the predicted contact probability matrix for each target and analyzed its relationship with the Top-5 precision of HF-S1 using 25 samples (Fig. 3a). Targets were stratified into high, medium, and low groups, revealing a strong correlation across datasets: lower-probability targets consistently yielded poorer predictions, while higher-probability targets were predicted with high precision. This indicates that low contact probability generally corresponds to more difficult targets, which often require additional sampling to achieve accurate predictions. When grouped by complex type (Fig. 3b), a similar trend is observed: Protein–protein targets typically show higher target-level contact probabilities, while protein–antibody complexes cluster in the lower range, indicating greater structural uncertainty. Protein–ligand complexes consistently exhibit high contact probabilities, suggesting that HF-S1 can often localize ligand binding sites with high confidence. In contrast, protein–RNA and protein–DNA complexes display a broader distribution, reflecting greater variability in prediction difficulty across these categories.